Chapter 20, Problem 026 In a hypothetical nuclear fusion reactor, the fuel is deuterium gas at a temperature of about 5.7 x 108 K. If this gas could be used to operate a Carnot engine with TL 817oC, what would be the engines efficiency? Number The number of significant digits is set to 7; the tolerance is +/-2% Click if you would like to Show Work for this question: Open Show Work Units
Answer #1

Th = 5.7*10^8 K

TL = 817+273 = 1090 K

engine efficiency is given by

n = 1 - TL / Th

n = 1 - 1090 / (5.7*10^8)

n = 0.9999981

or

n = 99.9981 %

============

do comment in case any doubt, will reply for sure.. Goodluck